OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36?

OpenStudy (anonymous):

y=6-x/2

OpenStudy (saifoo.khan):

For slope, solve for y. y = mx + b "m" is the slope.

OpenStudy (saifoo.khan):

For y-int make x = 0

OpenStudy (nenadmatematika):

-1/2 is the slope and y intercept is the point (0,6)

OpenStudy (anonymous):

y-intercept: (0, 6) x + 2y = 12 => y = -x/2 + 6

OpenStudy (anonymous):

slope is -1/2